# Quasi-Clique Discovery via Energy Diffusion

### v2

@@ 4.7 Empirical Validation of the Energy-Density Premise
4.8 Parameter Sensitivity
This section provides the detailed results of our parameter sensitivity analysis, which was designed to evaluate the impact of the number of diffusion rounds $T$ and the activation threshold $\theta$ on EDQC’s performance. For this study, we selected three representative datasets to cover different scales: ego-facebook (a small graph, $|V|\leq 10^{4}$), rgg_n_2_19_s0 (a medium graph, $10^{4}<|V|\leq 10^{6}$), and web-Google (a large graph, $|V|>10^{6}$).
On each of these graphs, we evaluated EDQC by testing 15 combinations of $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$. This evaluation was repeated under three different density thresholds, $\gamma\in\{0.7,0.8,0.9\}$, to ensure a comprehensive analysis. We report both the discovered quasi-clique size and the runtime for each configuration. The results for the small, medium, and large graphs are presented in Figure 57.
@@ 4.8 Parameter Sensitivity
Figure 5: Parameter sensitivity analysis for EDQC on the small representative graph (ego-facebook). Each of the three rows corresponds to a different density threshold $\gamma$. For each $\gamma$, the left plot shows the resulting quasi-clique size versus the activation threshold $\theta$ (on a log scale), and the right plot shows the corresponding runtime.
Figure 6: Parameter sensitivity analysis on the medium representative graph (rgg_n_2_19_s0). The layout of the plots is identical to that described in the caption of Figure 5.
Figure 7: Parameter sensitivity analysis on the large representative graph (web-Google). The layout of the plots is identical to that described in the caption of Figure 5.
Across all three datasets, a consistent pattern emerges. In terms of solution quality (size), configurations with $T=2$ and $T=3$ consistently and significantly outperform $T=1$, especially for smaller values of the activation threshold ($\theta\leq 10^{-3}$). As $\theta$ increases, making vertex activation stricter, the performance of all settings tends to decline. Regarding efficiency, the runtime plots clearly show that the computational cost increases with $T$. While $T=3$ occasionally finds a marginally larger quasi-clique than $T=2$, its runtime is substantially higher across all settings. For instance, in the large graph setting (Figure 7), the runtime for $T=3$ can be several times that of $T=2$ for little to no improvement in solution size.
Based on this analysis, we conclude that the configuration $T=2$ provides the best trade-off between solution quality and computational cost. Paired with a small activation threshold such as $\theta=10^{-4}$, which consistently lies in the stable, high-performance region shown in the plots, this configuration represents an effective and efficient choice. This justifies our use of $T=2$ and $\theta=10^{-4}$ for all main experiments presented in the paper.
5 Conclusion
In this paper, we introduced EDQC, a novel algorithm for quasi-clique discovery designed to mitigate key limitations of existing methods, such as their sensitivity to random seeds and lack of explicit edge-density guarantees. We demonstrated that the effectiveness of EDQC stems from its two-stage design. First, an adaptive energy diffusion process generates an energy ranking that highlights structurally cohesive regions. Then, guided by this ranking, a principled extraction mechanism based on conductance minimization identifies and refines a high-quality subgraph. To the best of our knowledge, EDQC is the first method to incorporate energy diffusion into this task, offering a robust alternative to traditional heuristics.
Our extensive experiments on 75 real-world graphs confirmed the effectiveness of this approach. The results demonstrated that EDQC finds larger quasi-cliques on most datasets, with consistently lower variance across runs, highlighting its superior stability. Furthermore, statistical tests confirmed the significance of these findings, and our analysis showed that EDQC maintains a competitive runtime, striking an excellent balance between solution quality and efficiency.
Future work can extend the EDQC framework in several promising directions. Generalizing the diffusion model for weighted and attributed graphs is an important next step. Applying EDQC to downstream tasks, such as anomaly detection in financial networks and community detection in social graphs, could further validate its practical utility. Finally, exploring more advanced refinement strategies may lead to further improvements in solution quality.

### v3

@@ 4.9. Empirical Validation of the Energy-Density Premise
4.10. Parameter Sensitivity
We studied the sensitivity of EDQC to the diffusion rounds $T$ and the activation threshold $\theta$ on three datasets of different scales (ego-facebook, rgg_n_2_19_s0, and web-Google). We tested $T\in\{1,2,3\}$ and $\theta\in\{10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2}\}$ under three values of $\gamma$. Due to space limits, Figure 6 reports results on web-Google.
Overall, EDQC is stable over a broad range of $(T,\theta)$. Increasing $T$ from $1$ to $2$ typically improves solution quality, while $T=3$ offers little additional gain but noticeably increases runtime. Smaller thresholds ($10^{-4}$ to $10^{-3}$) yield similar solution quality, whereas larger values ($\theta\geq 5\times 10^{-3}$) tend to over-prune and slightly degrade performance, especially at higher $\gamma$.
@@ 4.10. Parameter Sensitivity
Figure 6. Sensitivity of EDQC to $(T,\theta)$ on web-Google under three density thresholds $\gamma$. Left: quasi-clique size; right: runtime (log-scale $\theta$).
5. Conclusion
This paper proposes EDQC, an effective framework for cohesive group discovery under explicit edge-density constraints. EDQC leverages a lightweight adaptive energy diffusion process to rank vertices and localize candidate regions in interaction graphs; guided by this ranking, it performs conductance-based extraction and density-feasible refinement to obtain a $\gamma$-quasi-clique that satisfies the target threshold. This design aims to alleviate practical limitations observed in existing heuristics, including sensitivity to randomness and the lack of explicit density guarantees.
Extensive experiments on 75 real-world graphs across varying density thresholds show that EDQC identifies larger mean $\gamma$-quasi-cliques in the vast majority of cases, with consistently lower variance across runs while maintaining competitive runtime, and statistical tests support the significance of these improvements. Overall, EDQC offers a robust primitive for cohesive group discovery in interaction graphs, aligning with graph-based recommender systems where explicitly constrained cohesive groups underpin tasks such as social recommendation, bundle discovery, and community-aware modeling. Future work includes extending EDQC to weighted and attributed graphs and integrating it into end-to-end recommender pipelines to quantify task-level impact.